5/ 10 pts Partial Question 3 A random sample of 22 fifth grade pupils have a grade point average as follows: 5, 5.5,4.5, 5, 5* Please notice that one of the answer is incorrect
Answer #1

\bar x = 4.977

s = 0.523

At alpha = 0.05, the critical values are t0.025, 21 = +/- 2.080

Reject H0, if t < -2.08 or, t > 2.08

t < -2.08

or, (\bar x - \mu)/(s/\sqrt n) < -2.08

or, (\bar x - 4.7)/(0.523/\sqrt 22) < -2.08

or, \bar x < -2.08 * (0.523/\sqrt 22) + 4.7

or, \bar x < 4.4681

t > 2.08

or, (\bar x - \mu)/(s/\sqrt n) > 2.08

or, (\bar x - 4.7)/(0.523/\sqrt 22) > 2.08

or, \bar x > 2.08 * (0.523/\sqrt 22) + 4.7

or, \bar x > 4.9319

Reject the null hypothesis at significance level 0.05 if \bar x > 4.9319 or if \bar x < 4.4681.